# Consumption Smoothing in Metropolis: Evidence from the Working-class Households in Prewar Tokyo

### 26 Jan 2024 v3

@@ B.5 Comparing Distributions of the Household Size
Figure B.5 illustrates the household size distribution measured in the census and THBS sample. The THBS does not include households with one people and with a large number of people (i.e., 9+). The number of households in the former class is $362$, accounting for only $6$% of the total number of households. The latter class has $294$ households, accounting for a smaller share with $5$%. Thus, the THBS sample covers most of the classes in the household sizes of Tsukishima. Although the THBS sample has a higher share of the category with four people than the census, the two-sample Kolmogrov-Smirnov test does not reject the null hypothesis of the equality of distributions with an asymptotic $p$-value of 0.759.808080This result is based on the test using all categories. The same test excluding the categories for one, nine, and $10+$ person does not reject the null with an asymptotic $p$-value of 0.938. This shows that the THBS sample has a similar household size distribution to the population of Tsukishima (Section 3.1). It is, however, essential to remember that the THBS sample focuses on couples and households with children. This current study does not consider the consumption behaviors of the single factory workers’ households.
@@
Notes: This figure shows the share of the household in each household size category to the total number of households in Tsukishima measured in the 1920 Population Census and the THBS samples. The two-sample Kolmogrov-Smirnov test does not reject the null hypothesis of the equality of distributions with an asymptotic $p$-value of 0.759. Sources: Created by the author using the Tokyo City Office (1922a, pp. 262–283) and the THBS dataset.
@@ B.6 Estimating Average Head’s Monthly Income
Comprehensive estimation effort on the time series figures of average wages in prewar Japan has been done in the Long-Term Economic Statistics. Specifically, the LTES estimates the average national salary series of “general occupations,” “manufacturing sector,” and “manufacturing occupations (by medium classification).” For the manufacturing occupations category, the average wage per day of factory workers in Tokyo City and Osaka city from 1917 to 1922 are reported (Ohkawa et al. 1967, p. 255). These values were estimated using the “Annual Statistics of the City of Tokyo” (hereafter, the ASCT) and “Osaka City Statistical Table.” However, the household survey includes the household head’s monthly income but no information on the daily wage. Hence, the “average daily wage per factory worker” estimated in the LTES cannot be directly compared to the monthly income of the head of household, which is obtained from the Tsukishima Survey. In this section, I use available wage and labor statistics to estimate the average monthly income of male factory worker households in Tokyo and compare it to the THBS heads’ average monthly earnings.
To clarify the problem, I estimate average monthly earnings most simply. Specifically, I use the ASCT (volume 17), which provides wage statistics as of the end of 1918, as a case study. The ASCT (Volume 17) provides the “average daily wage of male factory workers aged 15 and over,” “number of factory workers aged 15 and over,” and “average number of working days per year” by industrial sector for four categories of factories: factories with ten or fewer workers, and factories using or not using motive power. Using the “number of workers over 15 years old” obtained for each factory category cell in each industrial sector as weights, I can get, for example, the average wage per month for a worker in the machinery sector as:
@@
This value is estimated at $27.8$ yen. This is the most straightforward estimate using daily wages, the number of workers, and the number of working days, and thus has several significant problems that create a downward bias in the estimates.
@@
Notes: This figure shows the density of the THBS heads’ age calculated from the THBS daily panel dataset. The dashed line (red) indicates the average age of all the heads (33.3 years old). The dashed line (blue) shows the average age of the heads, excluding an elder head aged 67.6 (32.3 years old). The difference between the two average figures is statistically insignificant with $p=0.4380$. Source: Created by the author using the THBS dataset.
@@
The first issue is that the average wages of factory workers listed in the ASCT (volume 17) are summarized as “15 years old or older.” Figure B.6 shows the THBS heads’ age distribution. The mean value is $33.3$ years old, which is unchanged (i.e., $32.3$ years old) if I exclude a head aged more than $60$. Only nine heads are under 29 years old, and most are around 30 to early 40s. This means I have to refer to the average daily wage for the 30s to 40s. However, since the number of factory workers is concentrated in the 20s, the average wage reported in the ASCT shall be considerably lower than the average wage for factory workers around 30 years old. To correct for this downward bias, I need the wage profiles by age.818181 The ASCT (volume 18), which contains statistics for 1919, does not list such wages by age group but only average wages. As is clear, the same problem arises in this case.
Second, there is no consideration of ancillary pay such as bonuses and various allowances. As explained later, these “accessory wages” were not negligible at the time. Thus, estimation without considering ancillary pay would have a non-negligible downward bias. In the following, I summarize the procedure for correcting such downward bias.
@@ B.6.1 Daily Wage Distribution by Ages
Since the early wage surveys focused on average wages, official statistics in the 1910s do not include average wage statistics by age (bin). By the “Statistical Data Field Survey Law” (Law No. 52) in April 1922, average wages by age group were surveyed for the first time in the “Labor Statistics Field Survey” in October 1924 conducted by the Social Affairs Bureau of the Ministry of Home Affairs (Committee for the History of the Labor Movement 1964, p. 14). The official reports of this survey in Tokyo were summarized as the Labor Statistics Table (hereafter, LST) for Tokyo City and Suburbs and published by the Tokyo City Statistics Division in 1926.
Figure B.7: Daily Wage by Age Bins: 1924 Manufacturing Census
@@
Notes: This figure shows the daily wage distribution by age bins measured in the manufacturing census in 1924. The weighted average daily wage for 30-39-year-old bin is shown in red. The weighted averages for all and 14+ bins are shown in gray and blue lines, respectively. The fitted fractional polynomial models are shown in the dashed lines. For the 1918 price data, the fitted model is $\hat{y}^{\textit{Daily Wage}}=-10.01+4.15a^{1/2}-0.33a^{2}$ with $R^{2}=0.99$. For the 1918 price data, the fitted model is $\hat{y}^{\textit{Daily Wage}}=-13.18+5.47a^{1/2}-0.43a^{2}$ with $R^{2}=0.99$. Sources: Created by the author using the Tokyo City Statistics Division (1926a, pp. 16; 244-245). Data on the consumer price index for the cities are obtained from Ohkawa et al. (1967, p. 135).
Figure B.7: Daily Wage by Age Bins: 1924 Manufacturing Census
Figure B.7 depicts daily wages by age for male factory workers in Tokyo City’s machinery and equipment manufacturing industry, based on statistics obtained from the LST. There are 13 age bins, and the number of workers and average wages are listed for each bin. Figure B.7a shows the distribution of wages adjusted to 1918 prices using the consumer price index (CPI) for urban areas (Ohkawa et al. 1967, p. 135). The weighted average for ages 14 and older is $2.58$ yen, which is about the same as the weighted average for daily wages of $2.56$ yen. The $2.58$ yen is between the average values for the 20-24 age bin and the 25-29 age bin. Thus, it may represent the daily wage of workers around 25. This is consistent with the fact that workers are concentrated in the 20s. However, the weighted average for the 30-39 age bin is $3.02$yen, which deviates significantly from the weighted average. The distribution of wages at 1919 prices shows a similar result (Figure B.7b). Using weighted averages for earnings comparisons leads to the erroneous conclusion that the sample households are upwardly biased.
To address these issues, I use the following steps to estimate the distribution of wages of the factory workers in Tokyo City in 1918-1919, the observation period of the THBS sample. Since the same estimation method is used for both years, I use 1918 as an example below.
@@
The dashed curve in Figure B.7 is the predicted distribution in equation 7. The precise fitting suggests that the predicted distribution works properly, especially below 50-54 bin.
Next, I shift the predicted distribution based on the difference in the average daily wages in the ASCT and LST. The weighted average of the daily wage is $2.56$ yen for the LST, whereas that for the ASTC is $1.05$ yen, meaning that there is roughly $1.5$ yen difference. [^83] For the LST, I use the number of factory workers in each age bin as weight. The ASCT reports the average daily wage for four different categories of factories in each manufacturing sector: the factory scale (10+ or below ten workers) and engines (use or not). I calculated the weighted average of the daily wage using the number of factory workers in each age bin (12–14 and 15+) and in each category as weights. Thus, I shift the functional polynomial function by this amount of difference to yield the proxy distribution for 1918.848484This operation, therefore, assumes that the wage-by-age distributions in 1918 and 1924 are reasonably similar. In the LST (i.e., the statistic on which the projected distribution is based), the weighted average for those aged 14 and older was 3.25 yen. In comparison, the overall weighted average was 3.23 yen, a difference of only 0.02 yen. A similar trend can be observed in the statistics obtained from the ASCT: the weighted average for those aged 15 and older is 1.14, while the overall average is 1.13, a difference of only 0.01 yen. At least, in the light of the first moment, this confirms the possibility that the wage distribution in the ASCT is similar to that in the LST. Figure B.8a illustrates this operation: dotted curve is a fractional polynomial shifted from the original fractional polynomial shown as a short dashed curve.
@@
The long dashed curve in Figure B.8a examines this relationship by fitting another fractional polynomial model using the 12-14-year old bin figure of the ASCT and the predicted wages for 25–65 from the shifted fractional polynomial.858585 The fitted model is as follows: $\hat{y}^{\textit{Daily Wage}}=-1.19+0.14a^{1}-0.002a^{2},\quad R^{2}=0.98.$ The deviance is reported to be $-149$. I used the 25–65 age range for the prediction based on the age bins used in the LST. Let this curve a predicted wage distribution for the true distribution. The average daily wage for the 30–40-year-old bin in the shifted fractional polynomial is $1.55$ yen, which is similar to that for the same-age bin in the predicted wage distribution (i.e., $1.52$ yen). However, the wage gaps for the younger ages show more significant differences. As long as our interest is in the late 20s to the early 40s, therefore, it is plausible to employ the predicted wages for the wage comparison conducted in Section 3.
@@
Notes: The short dashed curve illustrates the fractional polynomial (FP) in Figure B.7. The dotted curve indicates the FP shifted by the difference in the average daily wage between the ASCT and LST. The ASCT documents the wage statistics for four different categories for each manufacturing sector: the scale of the factory (10+ or below ten workers) and engines (use or not). I calculated the weighted average of the daily wage using the number of factory workers in each age bin and the category as weights. The long-dashed curve shows the FP refitted using the average daily wage for 12–14-year-old bin ($0.42$ yen) measured in the ASCT and the daily wages for 25–65-year-old suggested in the shifted FP. The average daily wage for a 12–14-year-old bin is only available for the ASCT of 1918. Therefore, this figure for the 1919 price is calculated using the ratio between the average daily wage ($1.05$ yen) and that for 12–14 years old bin ($0.42$) in 1918 as $1.80\times(0.42/1.05)$, where $1.80$ is the average daily wage measured in the ASCT of 1919. The red dashed line is the average daily wage for 30–40-year-old in the shifted FP. The blue dashed line is the average daily wage for 30–40-year-old in the refitted FP. Sources: Created by the author using the Tokyo City Office (1921, pp. 726–757); Tokyo City Statistics Division (1926a, pp. 16; 244-245).
@@
As I explained later, the average size of factories in Tsukishima was larger than in the entire city (Panel C in Table B.3). Since the larger factories had at least larger ancillary wages, the factory workers in Tsukishima might have had more earnings than those in the entire city (Online Appendix B.6.2). Therefore, it is important to remember that the average daily wages calculated in Figure B.8 are conservative estimates for the average daily wage for the skilled workers in the machinery sector in Tsukishima. However, given the scarcity in the wage-by-age statistics, I refer to these wages in the monthly wage calculation in Online Appendix B.6.3.
@@ B.6.2 Ancillary Wage
The daily wage estimated in Section B.6.1 is a fixed salary and, thus, does not include additional salaries (bonuses and allowances). Some data on bonuses and allowances in prewar Japan are available for the pre-Showa period, but unfortunately, there are no systematic statistics (Committee for the History of the Labor Movement 1964, pp. 14–15). In this section, therefore, I use available data from a detailed survey of factory workers’ salaries to determine the percentage of wages outside of regular pay. In the following, wages other than regular salaries, such as bonuses and allowances, are collectively referred to as ancillary wages.
Since the Tsukishima Survey targets skilled worker households, knowing the ancillary wages paid to skilled workers is necessary. The “Wage Survey Report” by Kitazawa (1924) is helpful in this regard. This report includes the result of a survey on the wages for machine finishers (fitters) and turners at six machine factories in Tokyo from May–October 1921. This is particularly useful as it examines the wage structure of skilled factory workers.
@@
There are two methods of paying wages to workers in the machine industry: one is to use hourly wages to calculate fixed salaries, and the other is to combine hourly wages with piece-rate wages, with the former being the most common (Kitazawa 1924, p. 33). However, wages paid in both methods are divided into fixed salaries and other ancillary wages, and there is not much difference in essence (Kitazawa 1924, p. 39). Fixed wages are salaries added from a fixed hourly wage rate, while ancillary wages include bonuses and overtime pay, night shift pay, and labor premiums (profit sharing). In the case of a combination of hourly and piece-rate pay systems, the accessory wage includes a share of the piece-rate profit.
@@
Panel A in Table B.3 summarizes the survey results. The surveyed factories included 86, 157, 376, 623, 928, and 2,837 workers, ranging from small to large. The percentage of skilled workers (finishers and turners) in the surveyed factories ranged from $27$% to $77$%, and the portion of monthly ancillary wages was 12–53%. Figure B.9 is a linear regression fit using log-transformed data on the percentage of ancillary wages and the number of workers. There is a positive correlation between the factory size and ancillary wages. This is consistent with the historical fact that larger factories have more extensive benefits than smaller ones (Odaka 1999).
@@
Next, I estimate the percentage of ancillary wages at the Tsukishima factories. One of the most frequently used statistics on factories in Tokyo is the Handbook of Factories (kōjyō tsūran), which covers all the factories with ten or more workers as of January 1, 1920. This handbook is the most comprehensive source for ascertaining the number of factories and their sizes around 1919. This document does not cover the number of tiny working places with less than nine workers. Importantly, however, the purpose of Panel C in Table 2 is to compare the average monthly wage of the skilled factory workers employed in the Tsukishima factories and the THBS heads’ average monthly earnings. Therefore, the censoring at the left tail of the factory size distribution is manageable for this purpose.
Columns 1 and 2 in Panel B of Table B.3 summarize the distributions of the number of factories, and the number of workers in Tsukishima reported in the handbook. Column 3 summarizes the average of the ancillary wage estimates predicted in Figure B.9. The average ancillary wage at the Tsukishima factory is about 12%, which is close to the value for the entire city of Tokyo (9%).
Panel C in Table B.3 compares the size of factories obtained from the handbook with the size of factories where the THBS heads in the machine industry worked. There are 32 (out of 40) heads of household for which the name of the factory where they worked is known. For 31 of these heads, the number of workers can be obtained from the handbook. In addition, three others are self-employed, two of whom are blacksmiths. Since these three were self-employed owners of tiny factories, I assume the number of workers to be 10, based on the report (Department of Health, Ministry of the Interior 1923a, p. 412). Column 1 of Panel C summarizes the distribution of plant sizes worked by these 35 heads of households. Compared to the distribution in Tokyo City and Tsukishima (columns 2–3), the THBS heads worked at the larger-scale factories.
Column 4 uses the same method as Panel B to calculate the average share of the ancillary wage for the THBS heads. The averages for the classes ranging from 10 to 199 workers are very similar to those of the THBS heads, Tsukishima, and Tokyo City classes. Despite the differences in the factory sizes, this supports the evidence that the THBS heads who worked in the small and medium-sized factories earned similar wage levels to the entire Tsukishima and Tokyo City.
However, in the 200+ bin, the average value of THBS heads (57%) is larger than the average value in Tokyo and Tsukishima (columns 3–4 in Panel B). This is because 13 heads of households work at Ishikawajima Shipyard, which has over 3,000 workers. Since there is a variation in the monthly earnings among the THBS heads working at the Ishikawajima Shipyard, the average value of THBS heads in the 200+ bin does not always represent the share of the ancillary wage for these heads. However, the percentage of workers in large factories is relatively higher in the THBS heads than in Tokyo City and Tsukishima. This suggests that these heads got a higher percentage of ancillary wages than the average. Therefore, the average monthly earnings of the THBS heads may be more or less higher than the average monthly income of the skilled workers in the Tsukishima factories.

### 25 Feb 2024 v6

@@ B.5 Comparing Distributions of the Household Size
Figure B.5 illustrates the household size distributions measured in the census and THBS sample. Pearson chi-squared test does not reject the null hypothesis of the equality of distributions with $p$-value of $0.948$. This supports the evidence that the THBS sample has a similar household size distribution to the population of Tsukishima (Section 3.1). As described in Section 3.2, while the THBS sample does not cover households with one person and with a very large number of people, it covers most of the classes in the household sizes with reasonable approximation to the population distribution. [^87] The number of households with one people is $362$, accounting for only $6$% of the total number of households. Similarly, the number of households with nine or more people is $294$, accounting for only $5$% (Section 3.2).
@@
Notes: This figure shows the share of the household in each household size category to the total number of households in Tsukishima measured in the 1920 Population Census and the THBS samples. Pearson chi-squared test does not reject the null hypothesis of the equality of distributions with $p$-value of $0.948$. The share in each household size measured in the census is used as the theoretical probability in calculating $\chi^{2}$ test statistic (Pearson 1900). Sources: Created by the author using the Tokyo City Office (1922a, pp. 262–283) and the THBS dataset.
@@ B.6 Estimating Average Head’s Monthly Income
Comprehensive estimation effort on the time series figures of average wages in prewar Japan was done in the Long-Term Economic Statistics (hereafter LTES). Specifically, the LTES estimates the average national salary series of “general occupations,” “manufacturing sector,” and “manufacturing occupations (by medium classification).” For the manufacturing occupations category, the LTES reports the average wage per day of factory workers in Tokyo City and Osaka cities from 1917 to 1922 (Ohkawa et al. 1967, p. 255). These values were estimated using the “Annual Statistics of the City of Tokyo” (hereafter, the ASCT) and “Osaka City Statistical Table.” However, the household survey includes the household head’s monthly income but no information on the daily wage. Hence, the “average daily wage per factory worker” estimated in the LTES is not useful for comparing the heads’ average monthly income from the THBS. In this section, I use available wage and labor statistics to estimate the average monthly income of male factory worker households in Tokyo and compare it to the THBS heads’ average monthly earnings.
To clarify the problem, I first estimate average monthly earnings most simply. Specifically, I use the ASCT (volume 17), which provides wage statistics as of the end of 1918, as a case study. For each industrial sector, the ASCT (volume 17) provides the “average daily wage of male factory workers aged 15 and over,” “the number of male factory workers aged 15 and over,” and “average number of working days per year” for four factory categories: factories with ten or fewer workers, and factories using or not using motive power. Using the “number of workers over 15 years old” in each factory category cell ($c$) as weights, I can get the average wage per month for a worker, for example in the machinery sector, as:
@@
This value is estimated to be $27.8$ yen. This is the most straightforward estimate using daily wages, the number of workers, and the number of working days. Thus, it has a few critical issues that cause a downward bias in the estimate.
@@
Notes: This figure shows the density of the THBS heads’ age calculated from the THBS daily panel dataset. The dashed line (red) indicates the average age of all the heads (33.2 years old). The dashed line (blue) shows the average age of the heads, excluding an elder head aged 67.6 (32.1 years old). The difference between the two average figures is statistically insignificant with $p=0.4087$. Source: Created by the author using the THBS dataset.
@@
The first issue is that the wages of factory workers are aggregated in a “15 years old or older” category in the ASCT (volume 17). Figure B.6 shows the THBS heads’ age distribution. The sample mean is $33.2$ years old, which is unchanged if I exclude a head aged more than $60$ (i.e., $32.1$ years old). This means that I have to refer to the average daily wage for the 30s to 40s for comparison. However, since many workers are in their 20s in the distribution of entire workers, the average wage reported in the ASCT shall be considerably lower than the average wage for factory workers around 30 years old. Thus, I need the wage profiles by age to correct for this downward bias.888888 The ASCT (volume 18), which contains statistics for 1919, does not list such wages by age group but only average wages. As is clear, the same problem arises in this case.
Second, there is no consideration of ancillary pay such as bonuses and various allowances. As explained later, these wages shared a non-negligible proportion of the head’s earnings. Thus, estimation without considering ancillary wages would have a strong downward bias. In the following, I summarize the procedure for correcting these downward biases.
@@ B.6.1 Daily Wage Distribution by Ages
Since the early wage surveys focused on average wages, official statistics in the 1910s do not include average wage statistics by age (bin). By the “Statistical Data Field Survey Law” (Law No. 52) in April 1922, average wages by age group were surveyed for the first time in the “Labor Statistics Field Survey” in October 1924 conducted by the Social Affairs Bureau of the Ministry of Home Affairs (Committee for the History of the Labor Movement 1964, p. 14). The result of this survey in Tokyo was published as the “Labor Statistics Table for Tokyo City and Suburbs” (hereafter, LST) by the Tokyo City Statistics Division in 1926.
Figure B.7: Average Daily Wage by Age Bins
@@
Notes: This figure shows the average daily wage distribution by age bins based on the manufacturing census in 1924. Figures B.7a and B.7b show the distributions under 1918 and 1919 prices, respectively. The weighted average daily wage for a 30-39-year-old bin is shown in red. The weighted averages for all and 14+ bins are shown in gray and blue lines, respectively. The fitted fractional polynomial models are shown in the dashed lines. For the 1918 price data, the fitted model is $\hat{y}^{\textit{Daily Wage}}=-10.01+4.15a^{1/2}-0.33a^{2}$ with $R^{2}=0.99$. For the 1919 price data, the fitted model is $\hat{y}^{\textit{Daily Wage}}=-13.18+5.47a^{1/2}-0.43a^{2}$ with $R^{2}=0.99$. Sources: Created by the author using the Tokyo City Statistics Division (1926a, pp. 16; 244-245). Data on the consumer price index for the cities are obtained from Ohkawa et al. (1967, p. 135).
Figure B.7: Average Daily Wage by Age Bins
Based on the LST, Figure B.7 shows the average daily wages of male factory workers in Tokyo City’s machinery and equipment manufacturing industry by age group. In the LST, there are 13 age bins, and the number of workers and average wages are listed for each bin. Figure B.7a shows the distribution of wages adjusted to 1918 prices using the consumer price index (CPI) for urban areas (Ohkawa et al. 1967, p. 135). The weighted average for ages $14$ and older is $2.58$ yen, which is about the same as the weighted average for daily wages of $2.56$ yen. The $2.58$ yen is between the average values for the 20-24 age bin and the 25-29 age bin. Thus, it likely to represent the daily wage of workers around 25. This is consistent with the fact that workers are concentrated in the 20s. However, the weighted average for the 30-39 age bin is $3.02$ yen, which deviates significantly from $2.58$ yen. The distribution of wages at 1919 prices shows a similar result (Figure B.7b). Using weighted averages for earnings comparisons leads to the erroneous conclusion that the sample households are upwardly biased.
To address this issue, I use the following steps to estimate the distribution of wages of the factory workers in Tokyo City in 1918-1919, the observation period of the THBS sample. Since the same estimation method is applied to both years, I use 1918 as an example below.
@@
The dashed curve in Figure B.7 is the predicted distribution in equation 8. The fitting suggests that the predicted distribution works properly, especially below the 50-54 bin.
Next, I shift the predicted distribution based on the difference in the average daily wages in the ASCT and LST. The weighted average of the daily wage is $2.56$ yen for the LST, whereas that for the ASTC is $1.05$ yen, meaning that there is roughly $1.5$ yen difference. [^90] For the LST, I use the number of factory workers in each age bin as weight. The ASCT reports the average daily wage for four different categories of factories in each manufacturing sector: the factory scale (10+ or below ten workers) and engines (use or not). I calculated the weighted average of the daily wage using the number of factory workers in each age bin (12–14 and 15+) and in each category as weights. Thus, I shift the functional polynomial function by this amount of difference to yield the proxy distribution for 1918.919191This operation, therefore, assumes that the wage-by-age distributions in 1918 and 1924 are reasonably similar. In the LST (i.e., the statistic on which the projected distribution is based), the weighted average for those aged 14 and older was 3.25 yen. In comparison, the overall weighted average was 3.23 yen, a difference of only 0.02 yen. A similar trend can be observed in the statistics obtained from the ASCT: the weighted average for those aged 15 and older is 1.14, while the overall average is 1.13, a difference of only 0.01 yen. At least, in the light of the first moment, this confirms that the wage distribution in the ASCT is similar to that in the LST. Figure B.8a illustrates this operation: dotted curve is a fractional polynomial shifted from the original fractional polynomial shown as a short dashed curve.
@@
The long dashed curve in Figure B.8a examines this relationship by fitting another fractional polynomial model using the 12-14-year old bin figure of the ASCT and the predicted wages for 25–65 from the shifted fractional polynomial.929292 The fitted model is as follows: $\hat{y}^{\textit{Daily Wage}}=-1.19+0.14a^{1}-0.002a^{2},\quad R^{2}=0.98.$ The deviance is reported to be $-149$. I used the 25–65 age range for the prediction based on the age bins used in the LST. Let this curve predict wage distribution for the true distribution. The average daily wage for the 30–40-year-old bin in the shifted fractional polynomial is $1.55$ yen, which is similar to that for the same-age bin in the predicted wage distribution (i.e., $1.52$ yen). However, the wage gaps for the younger ages show more significant differences. As long as our interest is in the late 20s to the early 40s, therefore, it is plausible to employ the predicted wages for the wage comparison in Section 3.
@@
Notes: The short dashed curve illustrates the fractional polynomial (FP) in Figure B.7. The dotted curve indicates the FP shifted by the difference in the average daily wage between the ASCT and LST. The ASCT documents the wage statistics for four different categories for each manufacturing sector: the scale of the factory (10+ or below ten workers) and engines (use or not). I calculated the weighted average of the daily wage using the number of factory workers in each age bin and the category as weights. The long-dashed curve shows the FP refitted using the average daily wage for 12-14-year-old bin ($0.42$ yen) measured in the ASCT and the daily wages for 25-65-year-old suggested in the shifted FP. The average daily wage for a 12–14-year-old bin is only available for the ASCT of 1918. Therefore, this figure for the 1919 price is calculated using the ratio between the average daily wage ($1.05$ yen) and that for 12–14 years old bin ($0.42$) in 1918 as $1.80\times(0.42/1.05)$, where $1.80$ is the average daily wage measured in the ASCT of 1919. The red dashed line is the average daily wage for 30–40-year-old in the shifted FP. The blue dashed line is the average daily wage for 30–40-year-old in the refitted FP. Sources: Created by the author using the Tokyo City Office (1921, pp. 726–757); Tokyo City Statistics Division (1926a, pp. 16; 244-245).
@@ B.6.2 Ancillary Wage
The daily wage estimated in Section B.6.1 is a fixed salary and thus, does not include additional salaries such as bonuses and allowances (hereafter, wages other than regular salaries are collectively referred to as ancillary wages). However, there are no systematic statistics on these ancillary wages throughout the prewar period (Committee for the History of the Labor Movement 1964, pp. 14–15). Moreover, since the Tsukishima Survey targets skilled worker households, information on the ancillary wages paid to skilled workers is necessary. Fortunately, I found a survey report named “Wage Survey Report” edited by Kitazawa (1924). This report includes the result of a survey on the wages for machine finishers (fitters) and turners at six machine factories in Tokyo from May–October 1921. Thus, this is particularly useful material to know the wage structure of skilled factory workers.
@@
According to Kitazawa (1924), there are two methods of paying wages to workers in the machine industry: one is to use hourly wages to calculate fixed salaries, and the other is to combine hourly wages with piece-rate wages, with the former being the most common (p. 33). Despite this, wages paid in both methods are usually divided into fixed salaries and other ancillary wages, and there is not much difference in essence (p. 39). Fixed wages are salaries added from a fixed hourly wage rate, while ancillary wages include bonuses and overtime pay, night shift pay, and labor premiums (profit sharing). In the case of a combination of hourly and piece-rate pay systems, the accessory wage includes a share of the piece-rate profit.
@@
Panel A in Table B.3 summarizes the survey results. The surveyed factories included 86, 157, 376, 623, 928, and 2,837 workers, ranging from small to large plants. The percentage of skilled workers (finishers and turners) in the surveyed factories ranged from $27$% to $77$%, and the portion of monthly ancillary wages was 12–53%. Figure B.9 is a linear regression fit using log-transformed data on the percentage of ancillary wages and the number of workers. There is a positive correlation between the factory size and ancillary wages. This is consistent with the historical fact that larger factories have more extensive benefits than smaller ones (Odaka 1999).
@@
Next, I estimate the percentage of ancillary wages at the Tsukishima factories. One of the most frequently used statistics on factories in Tokyo is the Handbook of Factories (kōjyō tsūran), which covers all the factories with ten or more workers as of January 1, 1920. This handbook is the most comprehensive source for ascertaining the number of factories and their sizes around 1919. This document does not cover the number of tiny working places with less than nine workers. Importantly, however, the purpose of Panel C in Table 2 is to compare the average monthly wage of the skilled factory workers in Tsukishima and the THBS heads’ average monthly earnings. Since the THBS heads did not work in the tiny workplaces, the censoring at the left tail of the factory size distribution is unlikely to influence this comparison.
Panel B of Table B.3 summarizes the distributions of the number of factories, the total number of workers, and the estimated average share of ancillary wage in Tokyo City (Column 1) and Tsukishima (Column 2). The average ancillary wage at the Tsukishima factories is 12%, which is slightly larger than the value for the entire city of Tokyo (9%). The percentage shares of the number of factory workers by each bin are listed in Columns (1) and (2) of Panel C in Table B.3. As shown, Tsukishima factories tended to have larger factories than those of the entire city, which is reflected in the slightly larger estimate in the average ancillary wage for the Tsukishima factories. In Column (3) of Table B.3, the number of THBS heads whose occupation, factory name, and factory size are available are listed. While the percentage shares of the heads in each bin show a similar distribution to that of the Tsukishima factories, the share of middle (large) sized factories is larger (smaller) than the population. If interpreted literally, this tendency implies that the (unobservable) true average ancillary wage for the THBS heads may be slightly lower than the population mean. [^93] Note again that the purpose of Panel C in Table 2 is to compare the average monthly wage of the skilled factory workers employed in the Tsukishima factories to the THBS heads’ average monthly earnings. Therefore, the purpose herein is to estimate the average ancillary wage for the target population, say the Tsukishima factories but not to calculate the average ancillary wage for the THBS sample itself.